# Automating Governing Knowledge Commons and Contextual Integrity (GKC-CI) Privacy Policy Annotations with Large Language Models

### v2

@@ Abstract.
Identifying contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policy texts can facilitate normative privacy analysis. However, GKC-CI annotation has heretofore required manual or crowdsourced effort. This paper demonstrates that high-accuracy GKC-CI parameter annotation of privacy policies can be performed automatically using large language models. We fine-tune 50 open-source and proprietary models on 21,588 GKC-CI annotations from 16 ground truth privacy policies. Our best performing model has an accuracy of 90.65%, which is comparable to the accuracy of experts on the same task. We apply our best performing model to 456 privacy policies from a variety of online services, demonstrating the effectiveness of scaling GKC-CI annotation for privacy policy exploration and analysis. We publicly release our model training code, training and testing data, an annotation visualizer, and all annotated policies for future GKC-CI research.
@@ 1. Introduction
Furthermore, no previous work on automating privacy policy annotation with machine learning (Wilson et al., 2018; Harkous et al., 2018; Bannihatti Kumar et al., 2020; Andow et al., 2020, 2019; Zimmeck et al., n. d.; Story et al., 2019; Tang et al., 2023) has been based on CI or GKC-CI. Rather, this prior work has focused on different sets of annotation tags that are not useful to the GKC/CI scholarly community as they do not produce results relevant to GKC/CI analysis. This motivates the development of novel automated techniques for the GKC-CI annotation task.
In this paper, we train a variety of large language models (LLMs) to perform automated GKC-CI parameter annotations of privacy policies. Doing so demonstrates the feasibility of large-scale longitudinal and cross-industry analysis of privacy policies using the GKC-CI framework. Specifically, we train and evaluate 50 LLMs from five different model families, ranging from open-source to proprietary models. We perform two ablation studies with the LLMs, and briefly remark on some key lessons which we believe may be broadly applicable to other researchers working in this space. We then proceed to a nuanced discussion of the behavior of our best performing model, particularly with respect to the model’s errors.
@@
  • We demonstrate that accurate GKC-CI parameter annotations of privacy policies can be performed automatically by a fine-tuned Large-Language Model (LLM), substantially improving scalability and reducing expense compared to manual and crowdsourcing approaches. Further, through training data review, we find these annotations are as accurate as those obtained from humans.
  • We train and publicly release the data and scripts necessary to reproduce a large language model capable of performing automatic GKC-CI annotation of privacy policies. Our automated approach has an average per-policy annotation cost between \$0.23-0.44, depending on the length of the policy. [^1] The cost estimation range is the average cost of annotating privacy policies from both the 164 least popular and the 164 most popular websites, per their Tranco (Pochat et al., 2018) ranking, in the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We note that reproducing our model currently incurs a one-time training cost of \$195.
@@ 2.1. The Usable Privacy Project
This line of research, while independently impressive, is not relevant for the community of scholars using contextual integrity or governing knowledge commons to interpret privacy policies and other legal documents. The sets of annotation tags used for the Usable Privacy Project are not analogous to the 5-parameter information flow description fundamental to CI nor the institutional grammar fundamental to GKC. This means that a separate thread of research, described below, has been needed to develop privacy policy annotation approaches useful for GKC/CI analyses.
@@ 2.2. Manual CI Annotation
Unlike the previous work by Wilson et al. (Wilson et al., 2016a, b, 2018) and the Usable Privacy Project, our work is based on Shvartzshnaider et al.’s GKC-CI framework (Shvartzshnaider et al., 2022), which provides a theoretically grounded basis for identifying ambiguity and potential privacy-violating behavior. Our work is less focused on helping consumers understand privacy policies than (Wilson et al., 2016a, b, 2018) and is more focused on automatable, longitudinal, and cross-industry analysis of privacy policies. Our work is also novel in its use of large-language models (the current state-of-the-art in natural language processing) and achieves considerably better performance than the machine learning annotation in (Wilson et al., 2018).
In 2019, Shvartzshnaider et al. (Shvartzshnaider et al., 2019a) used the theory of contextual integrity (CI) (Nissenbaum, 2009) to inform a new approach to privacy policy annotation. This approach seeks to identify the five information flow parameters defined by CI (Section 3) in privacy policy text. CI parameter annotation enabled the identification of ambiguities in information transfer descriptions. Shvartzshnaider et al. (Shvartzshnaider et al., 2022) combined contextual integrity with governing knowledge commons (GKC) (Frischmann et al., 2014; Sanfilippo et al., 2018) in 2022 to create a combined GKC-CI framework. GKC-CI extends the potential scope of CI annotation to eight total parameters, four from CI and four from the GKC institutional grammar (Section 3).
@@
The scalability issues posed by crowdworker annotation clearly motivate this study, which seeks to automate CI parameter annotation through the use of large language models (LLMs). Our work is also novel in its use of the expanded eight-parameter GKC-CI labels as annotation tags (Section 3) rather than the five-parameter CI tags used in Shvartzshnaider et al.’s original paper.
We apply our automated method to annotate a large corpus of privacy policies, including up to 20 years of longitudinal policies from 8 major technology companies (128 policies) and 328 contemporary policies from across the technology industry. This is orders of magnitude more privacy policies than have been manually annotated for GKC-CI research in previous work (Shvartzshnaider et al., 2019a).
@@ 2.3. Privacy Policy Analysis With Machine Learning
Several other studies have also applied machine learning to privacy policies, although none have been based on nor produced output useful for GKC-CI research. In 2018, Harkous et al. (Harkous et al., 2018) trained a hierarchy of convolutional neural networks to build a Question-Answering system that supports free-form querying of privacy policy content. Other ML-based approaches essentially parse privacy policies for information of interest, such as by using a logistic regression model to identify opt-out statements in privacy policy text (Bannihatti Kumar et al., 2020). PoliCheck (Andow et al., 2020), an expansion of PolicyLint (Andow et al., 2019), is capable of differentiating between first-party and third-party entities in flow-to-policy consistency analysis. Zimmeck et al. (Zimmeck et al., n. d.) and Story et al. (Story et al., 2019) used support vector machines to identify non-compliance between Android application code and the applications’ privacy policies. Their approach could be used to highlight these statements for consumers to make opt-out decisions without needing to read the entire policy themselves. Our application of machine learning to GKC-CI privacy policy annotation is similarly tightly focused, but on a task that does not overlap these earlier works.
More recently, some research groups have utilized LLMs broadly in the legal space or more specifically for privacy policy analysis. LLM benchmarks in the legal space include that of (Dai et al., 2023) and (Fei et al., 2023). Further, Ravichander et al. (Ravichander et al., 2019) trained a BERT-based large language model to answer questions about privacy policies in a Q&A format (not annotation) using a corpus of 1750 questions and 3500 expert answers. Other work that adapts LLMs for legal question-answering include (wan et al., 2023) and (Yue et al., 2023).
However, others have used LLMs specifically for privacy policy analysis. Tang et al., used various LLMs to annotate privacy policies via prompting  (Tang et al., 2023). However, the annotations are very simple (e.g., ”1st Party Collection”) and could likely be found using regexes. In contrast, our annotation approach uses a nuanced and theoretically grounded framework (GKC-CI) for normative privacy analysis. We believe LLMs are particularly suited for nuanced annotations as opposed to classical neural-network methods. This is because of LLMs’ increased representational capacity as well as their training on internet webscrapes, which exposes LLMs to cultural norms (Vaswani et al., 2017; Radford and Narasimhan, 2018; Devlin et al., 2018). However, we cannot provide a comparison to other researcher’s classical approaches for automated GKC-CI annotation because there are none developed in this space. To address this, we train an RNN on this GKC-CI task in our work.
@@ 4. Methods
We describe how we train LLMs to perform GKC-CI parameter annotation for privacy policies as accurately as possible. In doing so, we contribute the following insights.
First, we remark on how different LLMs of various architectures and sizes effect performance through ablation experiments. Because there is no other related work demonstrating how NLP models may be applied for GKC-CI annotation, we believe these insights may be particularly helpful to those considering normative privacy policy analysis in this space. Details of the model training process are provided in Section 4.4 and performance details are provided in Section 5. After evaluating a broad selection of models to find the most suitable candidate, we further tune the hyperparameters of our best performing model.
Second, we qualitatively examine the errors made by our best performing model to gain deep insights over the model’s behavior, particularly as it relates to practical deployment. We similarly believe that our approach to error analysis may provide a useful foundation for how others may analyze a model’s behavior in this space. The results of our error analysis are reported in Section 5.4.
Finally, we again use our best performing model to annotate a large set of privacy policies (Amos et al., 2021), which we report in Section 7, providing examples of the types of longitudinal and cross-industry analyses that can be performed via automated GKC-CI annotation.
@@ 4.1. Training and Testing Data
Our ground-truth labels are obtained by manually annotating GKC-CI parameters in 16 privacy policies from popular online services and e-learning websites, the exact breakdown of which is shown in Table 3 of Appendix A. We downloaded these privacy policies in HTML format and converted them to plain text for annotation. We used a customized version of the Brat Rapid Annotation Tool (Stenetorp et al., 2012) to manually label all GKC-CI parameters in the policies. In order to achieve consistent annotations across all annotators, we used a fixed set of guidelines defining each of the GKC-CI parameters (Appendix B). These guidelines were taken from (Shvartzshnaider et al., 2019a) for CI parameters and (Shvartzshnaider et al., 2022) for GKC-CI parameters to ensure continuity with prior work. Our ground-truth annotations included 6781 GKC-CI parameters across all 16 policies (Table 3). This ground-truth annotation process took two research assistants one semester to perform, including time spent learning the task.
@@ 4.2. Formatting Examples
In formatting our examples, we wanted to ensure that the model is able to find the relevant text to be annotated given the presence of a parameter. Note that this necessitates teaching the model to filter through irrelevant text as well. As such, we include positive examples (examples where a parameter is present) and negative examples (text where a given parameter is not present, i.e., irrelevant text). The inclusion of negative examples is necessary to ensure that the model is usable in a real-world environment. Specifically, not every sentence of a privacy policy will include a GKC-CI parameter. By including negative examples, we ensure that our model will only output a parameter if one is present in its input sentence. We develop negative examples by taking positive examples and changing the parameter tag to one which does not occur in the sentence. We create a negative example for all possible negative tags. An example of a negative example is shown in the text immediately preceding Section  4.2.1.
We lightly format each sentence of a privacy policy as the basis of an example. We use sentences as the atomic unit for model input because sentence divisions are natural delimiters and previous work has shown reasonable accuracy with sentence-based annotation  (Wilson et al., 2018). We observe that other natural units of separation, such as paragraphs, tended to be particularly long due to the legal nature of privacy policies. This length ultimately resulted in problems where the paragraphs were too long to fit into the context window of some of our LLMs. [^4] Further, to our knowledge resolving a finite-length context window with large quantities of text is an open problem in the NLP space, although there has been substantial work in the area such as  (Press et al., 2022) and  (Beltagy et al., 2020). To ensure that training data was held consistent between LLMs, and because of the prior work above, we use sentences as the basis of our annotations. However, we also perform supplemental analysis with GPT-3.5 to verify that we are not critically disadvantaging our model by only feeding in sentences. The results of this are shown in Appendix  C.
Specifically, each example consists of the following parts: (1) a prefix to orient the LLM to the task, (2) a sentence from a privacy policy, (3) the GKC-CI parameter of interest, and (4) text delimiters. We include the text delimiters because modern LLMs decide what text to generate next based on all the text in their context window. As such, they cannot by default determine what text has been provided via the prompt, and what the LLM has generated. We choose the extremely minimal prefix “Annotate: ” for our task. We chose this minimal prefix to minimize the effects of prompt choice while still leveraging the training benefits of using a prompt  (Webson and Pavlick, 2022; Scao and Rush, 2021). We chose our text delimiters based on recommendations in OpenAI’s documentation, namely “–¿” and “x-x-x” respectively. Examples of how we fully format our examples are shown below:
@@ 4.2.1. OpenAI models
Some chatbot models under OpenAI take an additional prompt, a system message, as input. GPT-3.5 Turbo is one such model. For models which take a system message, we replace the default system message from “You are a helpful assistant.” to “You are an assistant that understands Helen Nissenbaum’s theory of Contextual integrity (CI) and the governance of knowledge commons framework (GKC). This framework is abbreviated as GKC-CI. You reply with brief, to-the-point answers with no elaboration.” We do so to help orient the model to our task.
Finally, we perform an experiment leveraging the fact that Turbo is designed to respond to conversational inputs, as it is a chatbot. Specifically, we change our prefix from Annotate: to “For the following excerpt, provide the GKC-CI annotation of ’¡parameter¿’: ”. We call the model produced under this intervention GPT-3.5 Turbo, Prompt Engineered. Note that this is distinct from our baseline models that we prompt without fine-tuning.

### v3

@@ Abstract.
Identifying contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policy texts can facilitate normative privacy analysis. However, GKC-CI annotation has heretofore required manual or crowdsourced effort. This paper demonstrates that high-accuracy GKC-CI parameter annotation of privacy policies can be performed automatically using large language models. We fine-tune 50 open-source and proprietary models on 21,588 ground truth GKC-CI annotations from 16 privacy policies. Our best performing model has an accuracy of 90.65%, which is comparable to the accuracy of experts on the same task. We apply our best performing model to 456 privacy policies from a variety of online services, demonstrating the effectiveness of scaling GKC-CI annotation for privacy policy exploration and analysis. We publicly release our model training code, training and testing data, an annotation visualizer, and all annotated policies for future GKC-CI research.
@@ 1. Introduction
Furthermore, no previous work on automating privacy policy annotation with machine learning (Wilson et al., 2018; Harkous et al., 2018; Bannihatti Kumar et al., 2020; Andow et al., 2020, 2019; Zimmeck et al., 2019; Story et al., 2019; Tang et al., 2023) has been based on CI or GKC-CI. Rather, this prior work has focused on different sets of annotation tags that are not useful to the GKC or CI scholarly communities as they do not produce results relevant to GKC or CI analysis. This motivates the development of novel automated techniques for the GKC-CI annotation task.
In this paper, we train a variety of large language models (LLMs) to perform automated GKC-CI parameter annotations of privacy policies. Doing so demonstrates the feasibility of large-scale longitudinal and cross-industry analysis of privacy policies using the GKC-CI framework. Specifically, we train and evaluate 50 LLMs from five different model families, ranging from open-source to proprietary models. We perform ablation studies with the LLMs, and briefly remark on some key lessons which we believe may be broadly applicable to other researchers working in this space. We then proceed to a nuanced discussion of the behavior of our best performing model, particularly with respect to the model’s errors.
@@
  • We demonstrate that accurate GKC-CI parameter annotations of privacy policies can be performed automatically by a fine-tuned large language model (LLM), substantially improving scalability and reducing expense compared to manual and crowdsourcing approaches. Further, through training data review, we find these annotations are as accurate as those obtained from humans.
  • We train and publicly release the data and scripts necessary to reproduce a large language model capable of performing automatic GKC-CI annotation of privacy policies. Our automated approach has an average per-policy annotation cost between \$0.23-0.44, depending on the length of the policy. [^1] This cost estimation range is the average cost of annotating privacy policies from both the 164 least popular and the 164 most popular websites, per their Tranco (Pochat et al., 2018) ranking, in the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We note that reproducing our model incurs a one-time training cost of \$195 at time of writing.
@@ 2.1. The Usable Privacy Project
This line of research, while independently impressive, is not relevant for the community of scholars using contextual integrity or governing knowledge commons to interpret privacy policies and other legal documents. The sets of annotation tags used for the Usable Privacy Project are not analogous to the 5-parameter information flow description fundamental to CI nor the institutional grammar fundamental to GKC. This means that a separate thread of research, described below, has been needed to develop privacy policy annotation approaches useful for GKC and CI analyses.
Unlike the work by Wilson et al. (Wilson et al., 2016a, b, 2018) and the Usable Privacy Project, our work is based on Shvartzshnaider et al.’s GKC-CI framework (Shvartzshnaider et al., 2022), which provides a theoretically grounded basis for identifying ambiguity and potential privacy-violating behavior. Our work is less focused on helping consumers understand privacy policies than (Wilson et al., 2016a, b, 2018) and is more focused on automatable, longitudinal, and cross-industry analysis of privacy policies. Our work is also novel in its use of large language models (the current state-of-the-art in natural language processing) and achieves considerably better performance than the machine learning annotation in (Wilson et al., 2018).
@@ 2.2. Manual CI Annotation
In 2019, Shvartzshnaider et al. (Shvartzshnaider et al., 2019a) used the theory of contextual integrity (CI) (Nissenbaum, 2009) to inform a new approach to privacy policy annotation. This approach seeks to identify the five information flow parameters defined by CI (Section 3) in privacy policy text. CI parameter annotation enables the identification of ambiguities in information transfer descriptions. In 2022, Shvartzshnaider et al. (Shvartzshnaider et al., 2022) combined contextual integrity with governing knowledge commons (GKC) (Frischmann et al., 2014; Sanfilippo et al., 2018) to create a combined GKC-CI framework. GKC-CI extends the potential scope of CI annotation to eight total parameters, four from CI and four from the GKC institutional grammar (Section 3).
@@
The scalability issues posed by crowdsourced annotation clearly motivate this study, which seeks to automate CI parameter annotation through the use of large language models (LLMs). Our work is also novel in its use of the expanded eight-parameter GKC-CI labels as annotation tags (Section 3) rather than the five-parameter CI tags used in Shvartzshnaider et al.’s original paper.
We apply our automated method to annotate a large corpus of privacy policies, including up to 20 years of longitudinal policies from 8 major technology companies (128 policies) and 328 contemporary policies from across the technology industry. This is orders of magnitude more privacy policies than have been manually annotated for CI research in previous work (Shvartzshnaider et al., 2019a).
@@ 2.3. Privacy Policy Analysis With Machine Learning
Several other studies have also applied machine learning to privacy policies, although none have been based on nor produced output useful for GKC-CI research. In 2018, Harkous et al. (Harkous et al., 2018) trained a hierarchy of convolutional neural networks to build a Question-Answering system that supports free-form querying of privacy policy content. Other ML-based approaches essentially parse privacy policies for information of interest, such as by using a logistic regression model to identify opt-out statements in privacy policy text (Bannihatti Kumar et al., 2020). PoliCheck (Andow et al., 2020), an expansion of PolicyLint (Andow et al., 2019), is capable of differentiating between first-party and third-party entities in flow-to-policy consistency analysis. Zimmeck et al. (Zimmeck et al., 2019) and Story et al. (Story et al., 2019) used support vector machines to identify non-compliance between Android application code and the applications’ privacy policies. Their approach could be used to highlight these statements for consumers to make opt-out decisions without needing to read the entire policy themselves. Our application of machine learning to GKC-CI privacy policy annotation is similarly tightly focused, but on a task that does not overlap these earlier works.
More recently, some research groups have utilized LLMs in the legal space broadly or for privacy policy analysis specifically. LLM benchmarks in the legal space include that of Dai et al. (Dai et al., 2023) and Fei et al. (Fei et al., 2023). Ravichander et al. (Ravichander et al., 2019) trained a BERT-based large language model to answer questions about privacy policies in a Q&A format (not annotation) using a corpus of 1750 questions and 3500 expert answers. Other work that adapts LLMs for legal question-answering include that of Wan et al. (Wan et al., 2023) and Yue et al. (Yue et al., 2023).
Tang et al. (Tang et al., 2023) used various LLMs to annotate privacy policies via prompting. However, the annotations are very simple (e.g., “1st Party Collection”) and could likely be found using regexes. In contrast, our annotation approach uses a nuanced and theoretically grounded framework (GKC-CI) for normative privacy analysis. We believe LLMs are particularly suited for nuanced annotations as opposed to classical neural network methods. This is because of LLMs’ increased representational capacity as well as their training on Internet webscrapes, which exposes LLMs to cultural norms (Vaswani et al., 2017; Radford and Narasimhan, 2018; Devlin et al., 2018). However, we cannot provide a comparison to other researcher’s classical approaches for automated GKC-CI annotation because there are none developed in this space. To address this, we train an RNN on the GKC-CI annotation task as a baseline for our work.
@@ 4. Methods
We next describe how we trained LLMs to accurately perform GKC-CI parameter annotation for privacy policies. In doing so, we make the following contributions.
First, we tested how various architectures and sizes of different LLMs effect performance through ablation experiments. After comparing a broad selection of models to find the most suitable candidate, we further tuned the hyperparameters of our best performing model. Details of the model training process are provided in Section 4, and performance results are provided in Section 5. Because there is no previous work demonstrating how NLP models may be applied for GKC-CI annotation, we believe that these results will be particularly helpful to those considering normative privacy policy analysis in this space.
Second, we qualitatively examined the errors made by our best performing model to gain insights into the model’s behavior, particularly as it relates to practical deployment. We believe that our approach to error analysis may provide a useful foundation for how others may analyze model behavior in this space. The details of our error analysis are reported in Section 5.4.
Finally, we used our best performing model to annotate a large set of privacy policies (Amos et al., 2021), which we report in Section 7, providing examples of the types of longitudinal and cross-industry analyses that can be performed via automated GKC-CI annotation.
@@ 4.1. Training and Testing Data
Our ground-truth labels were obtained by manually annotating GKC-CI parameters in 16 privacy policies from popular online services and e-learning websites, the exact breakdown of which is shown in Table 3 of Appendix A. We downloaded these privacy policies in HTML format and converted them to plain text for annotation. We used a customized version of the Brat Rapid Annotation Tool (Stenetorp et al., 2012) to manually label all GKC-CI parameters in the policies. In order to achieve consistent annotations across all annotators, we used a fixed set of guidelines defining each of the GKC-CI parameters (Appendix B). These guidelines were taken from (Shvartzshnaider et al., 2019a) for CI parameters and (Shvartzshnaider et al., 2022) for GKC-CI parameters to ensure continuity with prior work. Our ground-truth annotations included 6781 GKC-CI parameters across all 16 policies (Table 3). This ground-truth annotation process took two research assistants one semester to perform, including time spent learning the task.
@@ 4.2. Formatting Examples
We lightly formatted each sentence of our ground-truth annotated privacy policies as the basis of our training and testing examples. Each formatted example consisted of the following parts: (1) a prefix to orient the LLM to the task, (2) a sentence from a privacy policy, (3) the GKC-CI parameter of interest, and (4) text delimiters. We chose the extremely minimal prefix “Annotate:” to minimize the effects of prompt choice while still leveraging the training benefits of using a prompt  (Webson and Pavlick, 2022; Scao and Rush, 2021). We included the text delimiters because modern LLMs decide what text to generate next based on all the text in their context window. As such, they cannot by default determine what text has been provided via the prompt and what the LLM has generated. We chose our text delimiters based on recommendations in OpenAI’s documentation, namely “–¿” and “x-x-x” respectively. Two fully formatted examples are shown below:
  1. (1)
@@
In formatting these examples, we wanted to ensure that the models learned to find the relevant text to be annotated. This necessitated teaching the models to filter through irrelevant text. As such, we used positive examples (text where a parameter is present, such as example (1) above) and negative examples (text where a given parameter is not present, such as example (2) above). The inclusion of negative examples was necessary to ensure that the models are usable in a real-world environment; not every sentence of a privacy policy will include a GKC-CI parameter. By including negative examples, our models learned to only output a parameter if one is actually present in an input sentence. We created the negative examples by taking positive examples and changing the parameter of interest to one which does not occur in the sentence. We did this for each positive example and all possible parameters not already in that example. For instance, the negative example (2) above was created by switching the parameter of interest in the positive example (1) above from “recipient” (which should be labeled in the example text) to “aim” (which should not).
We used sentences as the atomic unit for model input because sentence divisions are natural delimiters and previous work has shown reasonable accuracy with sentence-based annotation  (Wilson et al., 2018). We observed that other natural units of separation, such as paragraphs, tended to be particularly long due to the legal nature of privacy policies. This length ultimately resulted in problems where the paragraphs were too long to fit into the context window of some of our LLMs. [^4] Further, to our knowledge, resolving a finite-length context window with large quantities of text is an open problem in the NLP space, although there has been substantial work in the area, such as  (Press et al., 2022) and  (Beltagy et al., 2020). To ensure that the training data was consistent between LLMs, and because of the prior work above, we used sentences as the basis of our annotations. However, we also performed a supplemental analysis with GPT-3.5 to verify that we were not critically disadvantaging our models by only feeding in sentences (Appendix C).
@@ 4.2.1. OpenAI Models
Some OpenAI chatbot models take an additional prompt, a system message, as input. GPT-3.5 Turbo is one such model. For these models, we replaced the default system message from “You are a helpful assistant.” to “You are an assistant that understands Helen Nissenbaum’s theory of Contextual integrity (CI) and the governance of knowledge commons framework (GKC). This framework is abbreviated as GKC-CI. You reply with brief, to-the-point answers with no elaboration.” We did so to help orient the model to our task.
Finally, we performed an experiment leveraging the fact that GPT-3.5 Turbo is designed to respond to conversational inputs, as it is a chatbot. Specifically, we changed the prefix from “Annotate:” to “For the following excerpt, provide the GKC-CI annotation of ‘¡parameter¿’: ”. We call the model produced under this intervention GPT-3.5 Turbo, Prompt Engineered. Note that this is distinct from our baseline models that we prompted without fine-tuning.